a card is picked at random from a standard deck of 52 playing cards. Find the odds that it is not a heart?
Answer by edjones(8007)

There are 13 hearts.
Probability that it is a heart is 13/52 = 1/4
Probability that it is not a heart is 39/52 = 3/4
Odds against it being a heart is 3/4 to 1/4 or 3 to 1
.
Ed